OpenStudy (anonymous):

Find the Derivative of f(x)=2x^2 -x-3. a=1

OpenStudy (anonymous):

Please show the step by step long method, not the power rule or whatever the shortcut is. They haven't taught us that yet, and we have to do it the tedious way for now. /

sam (.sam.):

\[f'(x)=\frac{f(x+\Delta x)-f(x)}{\Delta x}\]

sam (.sam.):

That's the first principle, substituting in \[f'(x)=\frac{2(x+\Delta x)^2-(x+\Delta x)-3-(2x^2-x-3)}{\Delta x}\] Simplify

OpenStudy (anonymous):

Yeah i did that. Except for Changeofx, i used h. Does it matter?

OpenStudy (anonymous):

I get 2x+1 or something else, but the correct answer is 4x-1.

sam (.sam.):

It doesn't

OpenStudy (anonymous):

Hence why i'm asking the question, to show all steps please.

OpenStudy (anonymous):

The correct answer is 4x-1. and the equation for the tangent line is y=3x-5. I don't see how.

OpenStudy (anonymous):

The final answers are shown for the first 2, and the one with the green star for the graph is correct. I'm asking how to obtain these.

sam (.sam.):

\[\frac{2(x+\Delta x)^2-x-\Delta x -3-2x^2 +x +3}{\Delta x}\]

OpenStudy (anonymous):

Yup already got that. thanks

sam (.sam.):

ok

OpenStudy (anonymous):

Except i expanded the first part, is that wrong?

OpenStudy (anonymous):

I have 2x^2 +2xh +h^2 then everything else you had.

OpenStudy (anonymous):

You don't have to take the time make your responses perfect, although i appreciate it. I'm more just looking for the explanation.

sam (.sam.):

\[\lim_{\Delta x \rightarrow 0}\frac{2(x^2+2x \Delta x + \Delta x^2)-2x^2-\Delta x}{\Delta x} \\ \\ \lim_{\Delta x \rightarrow 0} \frac{\cancel{2x^2}+4x \Delta x +2\Delta x ^2\cancel{-2x^2}-\Delta x}{\Delta x} \\ \\ \lim_{\Delta x \rightarrow 0} \frac{4x \Delta x+2\Delta x^2 -\Delta x}{\Delta x} \\ \\ \lim_{\Delta x \rightarrow 0}4x+2\Delta x +1\] Then apply limit you get 4x+1

OpenStudy (anonymous):

Oh i see, but wait should't it be 4x+2h-1. Not plus 1.

OpenStudy (anonymous):

The answer is 4x-1 anyways.

OpenStudy (anonymous):

Which then simplifies to 4x-1, not 4x+1.

sam (.sam.):

oh made a mistake there sorry, yeah its 4x-1

OpenStudy (anonymous):

no problem. Also why is the limit 0? I thought a=1?

OpenStudy (anonymous):

If you look at my screenshot, i don't see anywhere saying the limit is 0.

sam (.sam.):

That is when f'(x)=a and y will be f(a)

OpenStudy (anonymous):

Whatever it works so i get it. Can you explain to me part B? what y= for the equation of the line tangent to the graph.

sam (.sam.):

You gotta find the gradient first, we know that \[f(x)=2x^2-x-3 \\ \\ f'(x)=4x-1\] and the turning point is when f(1) \[f'(1)=4-1=3\]

OpenStudy (anonymous):

Okay, and then i use the forumla y-1=3(something goes here not sure what) Right?

sam (.sam.):

Then, from the given turning point, (a,f(a)), the y is f(a), so back to the original equation, \[y=f(a)=2a^2-a-3\] when a=1, \[f(1)=-2\] So this is your y

sam (.sam.):

Then use the point slope formula, \[y+2=3(x-1)\]

sam (.sam.):

Where coordinates are (1,-2) and m=3

OpenStudy (anonymous):

Oh okay i think i got it.

sam (.sam.):

good luck :)

OpenStudy (anonymous):

I tried it on another problem, and i can't seem to get it.

OpenStudy (anonymous):

f(x)=2x^2 +3x-2 when a=-2

OpenStudy (anonymous):

The derivative is 4x+3, i got that right. But i can't figure out the y=. If i plug in -2 for a back into the original equation, it comes out to 0.

OpenStudy (anonymous):

Which leaves me with y=3(x-2) which is 3x-6. But it is incorrect? :

sam (.sam.):

you have the question paper?

OpenStudy (anonymous):

Its the same as the last one.

OpenStudy (anonymous):

Except f(x) = 2x^2 +3x-2. when a=-2

OpenStudy (anonymous):

The derivative is 4x+3 i got it right. I just don't know how to solve for y the tangent line thing.

OpenStudy (anonymous):

Turns out its y=-5x-10. flutter got it wrong again, gotta do the whole problem over again. What did i do wrong.

sam (.sam.):

Is the actual answer y=-5x-10?

OpenStudy (anonymous):

Stupid mistake, i think i get it now. I re read what you wrote, was a little confusing but figured it out.

sam (.sam.):

okay

OpenStudy (anonymous):

Thanks

sam (.sam.):

yw